OpenStudy (whalenjacob):

One real-world problem that is solved by an equation Kevin had $15 with him. His mother gave him $10 and his father also gave him some money. After this, he had a total of $37 with him. How much money did his father give him?

OpenStudy (whalenjacob):

i just need the work not the answer

OpenStudy (oraclethinktank):

Okay, Kevin had $15 on him, right?

OpenStudy (whalenjacob):

yesh

OpenStudy (oraclethinktank):

We will use a b and c a=15, b= 10, and c= x

OpenStudy (whalenjacob):

ok

OpenStudy (oraclethinktank):

We know Kevin has 15 and his mom gave him 10, so we have 15 + 10 + x =?

OpenStudy (whalenjacob):

25

OpenStudy (oraclethinktank):

No.

OpenStudy (oraclethinktank):

BEcause it's 15 + 10, PLUS x

OpenStudy (oraclethinktank):

So for now it's undefined. We know that the total is 37.

OpenStudy (oraclethinktank):

The equation is now 15 + 10 + x =37.

OpenStudy (whalenjacob):

oh duh

OpenStudy (oraclethinktank):

So 15 + 10 =?

OpenStudy (whalenjacob):

ok

OpenStudy (whalenjacob):

x

OpenStudy (oraclethinktank):

No. x is the missing number Remember, a= 15, b = 10, c= x

OpenStudy (whalenjacob):

so is it c? idk im confused

OpenStudy (oraclethinktank):

No. a is what Kevin has, b is what his mom gave him, and c is what his dad gave him.

OpenStudy (oraclethinktank):

We know a and b, but not c

OpenStudy (oraclethinktank):

We also know that if you add a b and c you get 37.

OpenStudy (oraclethinktank):

So you add a and b which is 15 + 25.

OpenStudy (oraclethinktank):

*15 + 10 rather.

OpenStudy (oraclethinktank):

So what is the result? of 15 + 10?

OpenStudy (whalenjacob):

25

OpenStudy (oraclethinktank):

Okay, so we now have (15+10) + x = 37

OpenStudy (oraclethinktank):

Solve the equation in the parentheses and you get 25 + x =37

OpenStudy (oraclethinktank):

Now, it's as simple and 37 -25

OpenStudy (whalenjacob):

12

OpenStudy (oraclethinktank):

We know the total is 37, and that Kevin has 25. meaning if you take 25 from 37, you take what he had and what his mom gave him, leaving only what his dad gave him.

OpenStudy (oraclethinktank):

So 15 + 10 + 12=?

OpenStudy (oraclethinktank):

Since we now know a= 15, b= 10, c=12

OpenStudy (whalenjacob):

37

OpenStudy (whalenjacob):

thank you so muc

OpenStudy (oraclethinktank):

No problem.